OpenStudy (anonymous):

2x-3/4y=-7:x+1/2y=0 elimination, at least x or y solved.

OpenStudy (whpalmer4):

Is that (3/4)*y and (1/2)*x or 3/(4y) and 1/(2x)?

OpenStudy (anonymous):

uh. 3/4. *y. and 1/2y. 1/2 is just y.

OpenStudy (anonymous):

?

OpenStudy (tkhunny):

@Fireflame13 Think on this a little. Printing in a book is not the same as writing in a line. You MUST remember your Order of Operations. Use parentheses to enforce it where it is ambiguous or difficult. (1/2)y = \(\dfrac{1}{2}y\) and is unambiguous and not confusing. 1/(2y) = \(\dfrac{1}{2y}\), and is pretty clear. 1/2y = \(\dfrac{1}{2}y\), but we really can't tell if you MEANT to write that and many will be confused by it.

OpenStudy (anonymous):

ok...

OpenStudy (tkhunny):

Now to the problem... 2x-(3/4)y=-7 x+(1/2)y=0 Multiply the second equation by -2 and see if anything presents itself.

OpenStudy (anonymous):

-2x-1=0

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

gosh!!!!

OpenStudy (tkhunny):

Did you add the equations? Something magic may occur.